# On pro-cdh descent on derived schemes

### v2

@@ Abstract.
We prove a ‘pro-cdh descent’ result for suitably connective localizing invariants and the cotangent complex on arbitrary qcqs derived schemes. As an application, we deduce that the $K$-groups of arbitrary qcqs schemes vanish below the negative of their valuative dimension.
@@ 1. Introduction
The algebraic $K$-groups of a blowup $\operatorname{Bl}_{Z}(X)$ of a quasi-compact, quasi-separated (qcqs, for short) scheme $X$ in a regularly immersed center $Z$ have been computed by Thomason [Tho93] (see also [CHSW08, Prop. 1.5]). In particular, the blowup square
${E}$${\operatorname{Bl}_{Z}(X)}$${Z}$${X}$
gives rise to a cartesian square of algebraic $K$-theory spectra and hence to a long exact sequence of algebraic $K$-groups. If the closed immersion $Z\hookrightarrow X$ is not regular, this will fail in general. However, if $X$ is Noetherian, and one takes infinitesimal information into account, one still gets a cartesian square. More precisely, if $Z(n)$ and $E(n)$ denote the $n$-th infinitesimal thickenings of $Z$ in $X$ and $E$ in $\operatorname{Bl}_{Z}(X)$, respectively, the square of pro-spectra
@@
is weakly cartesian and hence induces a long exact sequence of pro-abelian groups. Note here that the pro-systems only depend on the underlying topological space of $Z$ (and $E$) as a closed subset of $X$ (and $\operatorname{Bl}_{Z}(X)$, respectively) but not on the particular subscheme structure. In fact, more generally any abstract blowup square of Noetherian schemes gives rise to a weakly cartesian square of $K$-theory pro-spectra. In this form, this result, often coined “pro-cdh descent”, [^1] We remark that it would be more appropriate to call this “pro-cdh excision” since it is not precisely a descent statement for some topology. However, together with Nisnevich excision, it does imply actual (Čech) descent for the pro-cdh topology of [KS24], see Theorem 6.1 there. For this reason, and to be in line with the existing literature we stick to the term “pro-cdh descent.” is first proven in [KST18], but there are important precursors [KS02, Cor06, Kri10, GH06, GH11, Mor16b, Mor18]. A variant for Noetherian ANS stacks has been considered in [BKRS22]. These pro-cdh descent results play a central role in the resolution of Weibel’s $K$-dimension conjecture [KST18] (and [BKRS22] for stacks) and the development of a continuous $K$-theory of rigid spaces [Mor16a, KST19, KST23].
Given the rising interest in non-Noetherian schemes, which appear naturally, for example, when working over a perfectoid base ring such as $\mathcal{O}_{\mathbb{C}_{p}}$, it is an obvious question, whether, or in which form, pro-cdh descent also holds in this general setting. The following example, a variant of which was constructed by Dahlhausen and the third author [DT22] precisely for that purpose and which was independently studied by the first two authors [Kel24, Footnote 2] with respect to the pro-cdh topology, shows that pro-cdh descent as formulated above does not hold for general abstract blowup squares of non-Noetherian schemes.
@@
A posteriori, this is not surprising: A key point in [KST18] was to prove Thomason’s blowup formula for derived blowups in quasi-smooth centers together with the fact that for every closed subscheme of an affine Noetherian scheme there exists a quasi-smooth derived closed subscheme with the same underlying classical scheme. The condition that all (derived) rings and schemes appearing be Noetherian then allowed to pass from derived schemes back to classical schemes. For instance, for a Noetherian commutative ring $A$ and elements $f_{1},\dots,f_{r}$ in $A$ generating the ideal $I\subseteq A$, the pro-system of (Koszul type) derived quotients $\{A/\!\!/f_{1}^{n},\dots,f_{r}^{n}\}_{n}$ is equivalent to the pro-system of discrete rings $\{A/I^{n}\}_{n}$. This suggested that some variant of the above pro-cdh descent statement, where one takes “derived” infinitesimal thickenings everywhere, could still be true. This is precisely what we achieve in this paper. Our main $K$-theoretic result is the following.
@@
Let $f\colon Y\to X$ be a proper, locally almost finitely presented morphism of qcqs derived schemes which is an isomorphism outside a closed subset $Z\subseteq|X|$ whose open complement is quasi-compact. Denote by $X_{Z}^{\wedge}$ and $Y_{Z}^{\wedge}$ the formal completion of $X$ (respectively $Y$) along $Z$ (respectively $f^{-1}(Z)$) viewed as ind-derived schemes. Then the square of pro-spectra
@@
Let us point out that this result is not specific to $K$-theory. It holds more generally for every localising invariant which is $k$-connective in the sense of [LT19] for some integer $k$; see Theorem 4.6 for the general statement. For instance, it also applies to $\operatorname{THH}$, $\operatorname{TC}$, and rational negative cyclic homology.
@@ Acknowledgement
We thank Christian Dahlhausen for discussions about pro-cdh descent in the non-Noetherian setting which led to the variant of Example 1.1 and which was one of the starting points of this project. We are grateful to Ben Antieau for allowing us to include his pro-excision result for derived blowups, which is stronger than the one obtained in [KST18]. We thank Matthew Morrow for bringing our attention to the cotangent complex and further comments and Mauro Porta for related discussions. Finally we thank David Rydh for enlightening comments on our Sections 2 and 3. In particular, he proposed to formulate the more precise version of Theorem 3.2 including different finiteness hypotheses.

### v3

@@ Abstract.
Grothendieck’s formal functions theorem states that the coherent cohomology of a Noetherian scheme can be recovered from that of a blowup and the infinitesimal thickenings of the center and of the exceptional divisor of the blowup. In this article, we prove an analogous descent result, called “pro-cdh descent”, for certain cohomological invariants of arbitrary quasi-compact, quasi-separated derived schemes. Our results in particular apply to algebraic $K$-theory, topological Hochschild and cyclic homology, and the cotangent complex.
As an application, we deduce that $K_{n}(X)=0$ when $n<-d$ for quasi-compact, quasi-separated derived schemes $X$ of valuative dimension $d$. This generalises Weibel’s conjecture, which was originally stated for Noetherian (non-derived) $X$ of Krull dimension $d$, and proved in this form in 2018 by Kerz, Strunk, and the third author.
@@ 1. Introduction
An example of an abstract blowup square of schemes is a cartesian square
(1) ${E}$${Y}$${Z}$${X}$$\scriptstyle{p}$$\scriptstyle{i}$
where $Y=\operatorname{Bl}_{Z}(X)$ is the blowup of a Noetherian scheme $X$ in a closed subscheme $Z$. More generally, an abstract blowup square of schemes is a cartesian square (1) of (potentially non-Noetherian schemes) where $i$ is a finitely presented closed immersion and the morphism $p$ is finitely presented, proper and an isomorphism over the open complement of $Z$ in $X$. Coverings $\{Z\hookrightarrow X,Y\to X\}$ for all abstract blowup squares as above together with the Nisnevich coverings generate Suslin and Voevodsky’s cdh topology [SV00], which plays an important role in the study of motives. It is an interesting question how cohomological invariants behave under such abstract blowup squares.
Thomason [Tho93] (see also [CHSW08, Prop. 1.5]) proved that for $X$ quasi-compact, quasi-separated (qcqs, for short), $Z\hookrightarrow X$ a regular closed immersion, and $Y$ the blowup of $X$ in $Z$, the square (1) gives rise to a (homotopy) cartesian square of algebraic $K$-theory spectra and hence to a long exact sequence of algebraic $K$-groups
$\dots\longrightarrow K_{i}(X)\longrightarrow K_{i}(Z)\oplus K_{i}(Y)\longrightarrow K_{i}(E)\longrightarrow K_{i-1}(X)\longrightarrow\dots$
For general abstract blowup squares (1), however, this will fail. If $X$ is Noetherian, one can repair this defect by taking infinitesimal information into account. More precisely, if $Z(n)$ and $E(n)$ denote the $n$-th infinitesimal thickenings of $Z$ in $X$ and $E$ in $Y$, respectively, Theorem A of [KST18] states that the square of pro-spectra
@@
is cartesian in a certain weak sense (see Section 4 for a definition). In particular, it induces a long exact sequence of pro-abelian groups. This result can be viewed as a $K$-theoretic version of Grothendieck’s formal function’s theorem [Gro61, Thm. 4.1.5]. It had several important precursors [KS02, Cor06, Kri10, GH06, GH11, Mor16b, Mor18]. Because of the role of abstract blowup squares in the cdh topology, such a result is often called “pro-cdh descent.” [^1] We remark that it would be more appropriate to call this “pro-cdh excision” since it is not precisely a descent statement for some topology. However, together with Nisnevich excision, it does imply actual (Čech) descent for the pro-cdh topology of [KS24], see Theorem 6.1 there. For this reason, and to be in line with the existing literature we stick to the term “pro-cdh descent.”
This pro-cdh descent result plays a central role in the resolution [KST18] of Weibel’s $K$-dimension conjecture, which states that the $K$-groups of a Noetherian scheme $X$ vanish below the negative of its Krull dimension, and in the development of a continuous $K$-theory of rigid spaces [Mor16a, KST19, KST23].
Given the rising interest in non-Noetherian schemes, which appear naturally, for example, when working over a perfectoid base ring such as $\mathcal{O}_{\mathbb{C}_{p}}$, it is an obvious question, whether, or in which form, pro-cdh descent holds in this general setting. The following example, a variant of which was constructed by Dahlhausen and the third author [DT22] precisely for that purpose and which was independently studied by the first two authors [Kel24, Footnote 2] with respect to the pro-cdh topology, shows that pro-cdh descent as formulated above does not hold for general abstract blowup squares of non-Noetherian schemes.
@@
It turns out that the reason for this failure is that we took classical quotients when forming the infinitesimal thickenings which discards torsion information. More concretely, let $A$ be a ring and $f$ an element of $A$. Derived algebraic geometry allows us to form the derived quotients $A/\!\!/f^{n}$ which are certain simplicial commutative rings. Then $\pi_{0}(A/\!\!/f^{n})$ is the classical quotient $A/f^{n}$, whereas $\pi_{1}(A/\!\!/f^{n})\cong\ker(f^{n}\colon A\to A)$ still carries the information about the $f^{n}$-torsion. We can then consider the pro-system $\{A/\!\!/f^{n}\}_{n}$ as a (derived) formal completion of $A$. For a Noetherian ring $A$, this formal completion is in fact equivalent to the usual (pro-)completion $\{A/f^{n}\}_{n}$. Our main theorem shows that with this modification, replacing infinitesimal neighbourhoods by “derived infinitesimal neighbourhoods,” pro-cdh descent does in fact hold in general. It is natural to formulate it in the generality of derived schemes. For a precise definition and the involved finiteness conditions, we refer to Section 2.
@@
Let $f\colon Y\to X$ be a proper, locally almost finitely presented morphism of quasi-compact and quasi-separated (qcqs for short) derived schemes which is an isomorphism outside a closed subset $Z\subseteq|X|$ whose open complement is quasi-compact. Denote by $X_{Z}^{\wedge}$ and $Y_{Z}^{\wedge}$ the formal completion of $X$ (respectively $Y$) along $Z$ (respectively $f^{-1}(Z)$) viewed as ind-derived schemes. Then the square of pro-spectra
@@
This result is not specific to $K$-theory. It holds more generally for every localizing invariant which is $k$-connective in the sense of [LT19] for some integer $k$; see Theorem 4.6 for the general statement. For instance, it also applies to topological Hochschild homology $\operatorname{THH}$, topological cyclic homology $\operatorname{TC}$, and rational negative cyclic homology.
@@ Acknowledgement
We thank Christian Dahlhausen for discussions about pro-cdh descent in the non-Noetherian setting which led to the variant of Example 1.1 and which was one of the starting points of this project. We are grateful to Ben Antieau for allowing us to include his pro-excision result for derived blowups, which is stronger than the one obtained in [KST18]. We thank Matthew Morrow for bringing our attention to the cotangent complex and further comments and Mauro Porta for related discussions. We thank David Rydh for enlightening comments on our Sections 2 and 3. In particular, he proposed to formulate the more precise version of Theorem 3.2 including different finiteness hypotheses. Finally, we thank the referee for a careful reading of the manuscript and useful comments.